OpenStudy (anonymous):

HELP:find the first partial derivative of f(x,y) integral from x to y cos(t^9) dt

OpenStudy (anonymous):

Can you restate the question? You want to find the first order partial derivative of a function f(x,y) with respects to both x and y correct? What exactly is your function?

OpenStudy (anonymous):

\[f(x,y)= \int\limits_{y}^{x} \cos (t^9) dt\]

OpenStudy (anonymous):

this is what i got

OpenStudy (anonymous):

\[1/x^9 \sin(x^9)+1/y^9 \sin(y^9) +C\]

OpenStudy (anonymous):

not sure if its right though

OpenStudy (anonymous):

ohhh..

OpenStudy (anonymous):

i am suppose to integrate by parts right?

OpenStudy (anonymous):

1/9sinx^9-1/9sinx^9+c

OpenStudy (anonymous):

In said situation, given theorized, ideal Pythagorean standards, one could, when, concluded extremities parametize area, fortitude of x with respect to inifinite limit gives, under mathematical conditions, that, stated integral has an unreal answer if eliminated factors cancel in a forthright fashion. All is explained by science

OpenStudy (anonymous):

i totally spaced out on that

OpenStudy (anonymous):

1/9sinx^9-1/9siny^9+c

OpenStudy (anonymous):

Ok lets see, I believe your integral is right, although there's no need for the +c because you have definite bounds...

OpenStudy (anonymous):

yes you are right no need +c

OpenStudy (anonymous):

so we don't integrate this by u substitution?

OpenStudy (anonymous):

or some other integration method?

OpenStudy (anonymous):

no need substitution

OpenStudy (anonymous):

hmm.......... actually

OpenStudy (anonymous):

Definite C, when referring to tangency, is, without question, derivized by cascadation of three bounds in the three dimensional coordinate plane, as defined by Cartesian math as well as Mendel, as, with respect to z, a common factor is pulled, given that a trifactor of polynomic trigonometry is frozen in intervals of, cosine, when not shown to be aligned, yields a C of zero which can, ultimately, not need be shown unless graphing is necessary.

OpenStudy (anonymous):

So we got 1/x^9 sin(x9)+1/y^9 sin(y9) when evaluated between x and y.... This means our integral without evaluating it, and in terms of t (aka indefinite integral) was.... 1/t^9 cos(t^9)... if you derive that in terms of t you do not get the original equation...

OpenStudy (anonymous):

so what does that mean?

OpenStudy (anonymous):

You are either missing part of the problem or your teacher is an retriceole. You can't solve this easily, if at all, because you are integrating for cosine with a function inside and nothing outside to return upon integrating. Try using a trig identity if you can think of one, or www.wolframalpha.com/. Otherwise you can just say unable to solve because it seems unreasonable for you to be asked to solve that if your teacher hasn't shown you how.

OpenStudy (anonymous):

Hmmm i actually had a problem just like this on my vector calc test a few weeks ago... I didn't get the problem but i think it has to do with cos(x^2) being an odd function... I'm looking for some sort of trig identity or trick now because neither u sub or by parts will work for something like this, and wolfram gives me a rediculious answer

OpenStudy (anonymous):

yeah we haven't learned this and there isn't an example in a book

OpenStudy (anonymous):

http://how-to-spell-ridiculous.com/

OpenStudy (anonymous):

You are going to need an incomplete gamma or knowledge of the Heaviside to solve this.

OpenStudy (anonymous):

hahahaha thanks ScienceWiz

OpenStudy (anonymous):

<3

OpenStudy (anonymous):

So I think I know the answer... If we use the Fundamental Theorem of Calculus we can state that the Integral of your original function then derivative in respects to t is just the original function right?

OpenStudy (anonymous):

So if you take the Integral of your original function f(x) then evaluate with respects to x and y. You'll get some F(x)-F(y) right.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

where F(x)= integral of x

OpenStudy (anonymous):

then we take the partial with respects to x of F(x)-F(y). F(y) goes to 0 because there are no x terms, while F(x) goes to the original function just in terms of x

OpenStudy (anonymous):

the same applies for your partial derivative of y... So you should just get for your\[f _{x}=\cos(x^9)\] and\[f _{y}=\cos(y^9)\]

OpenStudy (anonymous):

Thoughts?

OpenStudy (anonymous):

yeah I am kind of confused because I am a bit rusty on the FTC

OpenStudy (anonymous):

I'm having a hard time following exactly what you are saying. The main issue that I see is there is no way to get around integrating for t at somepoint. And you can't do that at the level of knowledge you are at right now. Also this link says no. And if you delete "from x to y" you can look at the indefinite integral. http://www.wolframalpha.com/input/?i=integral%28cos%28t^9%29dt%29+from+x+to+y

OpenStudy (anonymous):

FTC just states The derivative of the Integral is your original equation..

OpenStudy (anonymous):

Ya but you can't apply that theorem to this situation unless you were to integrate again, which would cause the same problem.

OpenStudy (anonymous):

hmmm? the initial problem states "find the first partial derivative of f(x,y) integral from x to y cos(t^9) dt" doesn't that mean you want to find the partial derivatives? Or do you just want the integral?

OpenStudy (anonymous):

I need to find the first partial derivatives fx and fy

OpenStudy (anonymous):

hmmm.. ok let me try explaining my logic better... So FTC states...\[d/dx(\int\limits_{a}^{x}f(t)dt) :=F'(t)= f(x)\] this means after you take the integral if you take the derivative you get your original function... So our function is f(x,y)=∫yxcos(t9)dt\[f(x,y)=\int\limits_{y}^{x}\cos(t^9) dt= \int\limits_{y}^{x}f (t)dt\] We can just rewrite our\[ \int\limits_{y}^{x}f (t)dt= F(x)-F(y) \] where F(x) is equal to the integral of our original equation evaluated at x... So now we take out function F(x)-F(y) and take the partials in respects to x and y...\[f _{x}(F(x)-F(y))= f(x)\] and \[f _{y}(F(x)-F(y))= -f(y)\] This is because when we take the partial of F(y) with respects to x it goes to 0...

OpenStudy (anonymous):

so f(x) = cos (x ^9) and -f(y)= -cos(y^9)? because of FTC

OpenStudy (anonymous):

Pretty much.

OpenStudy (anonymous):

that seems much easier :) I appreciate your help!! THANKS